# Spectrum radii of trees

### v1

@@ Abstract
For any positive integer $r$ and positive number $\alpha$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set of positive numbers defined recursively: $\alpha\in{\mathscr{W}}_{r}(\alpha)$, and for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):1\leq i\leq s\}$, where $1\leq s<r$, $\beta:=\alpha-\sum_{i=1}^{s}q_{i}^{-1}$ belongs to ${\mathscr{W}}_{r}(\alpha)$ as long as $\beta>0$. We first show that there exists a tree $T$ such that its maximum degree $\Delta(T)$ is at most $r$ and its spectrum radius $\lambda(T)$ is equal to $\alpha$ if and only if $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$. It follows that the set of spectrum radii of non-trivial trees is exactly the set of positive numbers $\alpha$ such that $\alpha^{-1}\in{\mathscr{W}}_{\lfloor\alpha^{2}\rfloor}(\alpha)$. Applying this conclusion, we prove that for any positive integers $r$ and $k$, there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\sqrt{k}$ if and only if $\frac{1}{4}k+1<r\leq k$.
@@
Keywords: tree, maximum degree, characteristic polynomial; eigenvalue; spectrum radius
@@ 1 Introduction
All graphs considered in this article are simple and undirected. For a graph $G$, let $V(G)$, $E(G)$ and $\Delta(G)$ denote its vertex set, edge set and maximum degree, respectively. For any non-empty subset $S$ of $V(G)$, let $G[S]$ denote the subgraph of $G$ induced by $S$. For any $u\in V(G)$, let $N_{G}(u)$ (or simply $N(u)$) denote the set of neighbours of $u$ in $G$, and let $d_{G}(u)$ or (simply $d(u)$) be the size of $N_{G}(u)$, called the degree of $u$ in $G$. Let $L(G)$ denote the set of vertices $v\in V(G)$ with $d_{G}(v)=1$.
For any integer $k>0$, let $\llbracket k\rrbracket=\{1,2,\dots,k\}$. For a graph $G=(V,E)$, where $V=\{v_{i}:i\in\llbracket k\rrbracket\}$, its adjacency matrix $A(G)$ according to the vertex ordering $v_{1},v_{2},\ldots,v_{n}$ is defined to be the $0-1$ matrix $(a_{i,j})_{n\times n}$, where $a_{i,j}=1$ if and only if $v_{i}v_{j}\in E$. The characteristic polynomial of $G$, denoted by $P(G,\lambda)$, is defined to be the function $det(\lambda I_{n}-A(G))$, where $I_{n}$ is the identical matrix of size $n$. It is known that $P(G,\lambda)$ is independent of the vertex ordering $v_{1},v_{2},\dots,v_{n}$, and the roots of the equation $P(G,\lambda)=0$, called the eigenvalues of $G$, are real numbers (see [6]), and the largest eigenvalue of $G$, denoted by $\lambda(G)$, is called the spectral radius of $G$.
This spectral radius of graphs, which has been studied extensively (see a survey [17] due to Stevanovíc), plays a crucial role in spectral graph theory, with applications in spanning combinatorics, network theory, and even theoretical physics (see [1, 2, 3, 4, 8]). In this article, we focus on the study of spectral radii of trees. Let $\Lambda$ denote the set of $\lambda(T)$ for all non-trivial trees $T$ (i.e., trees of order at least $2$). It is known that $q\geq 1$ for each $q\in\Lambda$. The first aim in this article is to provide a characterization of the set $\Lambda$. For any positive number $\alpha$ and positive integer $r$, we define a set ${\mathscr{W}}_{r}(\alpha)$ of positive numbers as follows:
@@
    for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $s<r$, $\beta:=\alpha-\sum\limits_{1\leq i\leq s}q_{i}^{-1}$ belongs to ${\mathscr{W}}_{r}(\alpha)$ whenever $\beta>0$.
It can be verified that ${\mathscr{W}}_{1}(\alpha)=\{\alpha\}$ for all $\alpha>0$, ${\mathscr{W}}_{r}(1)=\{1\}$ for all $r\geq 2$, ${\mathscr{W}}_{2}(\sqrt{2})=\{\sqrt{2},(\sqrt{2})^{-1}\}$ and ${\mathscr{W}}_{2}(2)=\{\frac{i+1}{i}:i\geq 1\}$. In this article, we apply the set ${\mathscr{W}}_{r}(\alpha)$ to determine if there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ for any given pair of positive numbers $\alpha$ and $r$, where $r$ is an integer.
@@
For any positive number $\alpha$ and positive integer $r$, there exists a tree $T$ such that $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ if and only if $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$.
For any tree $T$, $|V(T)|\geq 3$ if and only if $\Delta(T)\geq 2$. In 2003, Stevanović [16] proved that for any tree $T$ with $|V(T)|\geq 3$,
$\frac{1}{4}\lambda(T)^{2}+1<\Delta(T)\leq\lambda(T)^{2}.$ (1)
By the definition of ${\mathscr{W}}_{r}(\alpha)$, we have ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r^{\prime}}(\alpha)$ whenever $r<r^{\prime}$. Thus, a characterization of $\Lambda$ follows from Theorem 1.1 and (1) directly.
@@
$\Lambda$ is exactly the set of positive numbers $\alpha$ such that $\alpha^{-1}\in{\mathscr{W}}_{\left\lfloor{\alpha^{2}}\right\rfloor}(\alpha)$.
By (1), $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$ is a necessary condition for the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$. We wonder if, for any $\alpha\in\Lambda$ and positive integer $r$, this condition is sufficient for the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$.
@@
Our second aim in this article is to apply Theorem 1.1 to prove Conjecture 1.3 for the case $\alpha=\sqrt{k}$, where $k\geq 2$ is an integer. Certainly, it includes the case $\alpha=k$ for any integer $k\geq 2$.
@@
For any positive integers $k$ and $r$, where $r\geq 2$, there exists a tree $T$ with $\lambda(T)=\sqrt{k}$ and $\Delta(T)=r$ if and only if $\frac{1}{4}k+1<r\leq k$.
The remainder of this article is structured as follows. In Section 2, we examine the set ${\mathscr{W}}_{r}(\alpha)$ and establish a connection between a tree $T$ with $\lambda(T)=\alpha$ and the elements of ${\mathscr{W}}_{r}(\alpha)$. In Section 3, we introduce a family of non-trivial directed trees $T$ with a unique source $z$ and a vertex-weight function $w$, ensuring that $w(v)\in{\mathscr{W}}_{r}(\alpha)$ for each $v\in V(T)\setminus\{z\}$, for any given positive numbers $r$ and $\alpha$, where $r$ is an integer. In Sections 4 and 5, we present proofs of Theorems 1.1 and 1.4, respectively. Finally, in Section 6, we discuss open problems arising from our work.
2 The set ${\mathscr{W}}_{r}(\alpha)$
Let $\alpha$ and $r$ be positive, where $r$ is an integer. By definition, ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r+1}(\alpha)$ and $0<q\leq\alpha$ for each $q\in{\mathscr{W}}_{r}(\alpha)$. The following result tells that when $r\leq\alpha^{2}/4+1$, all numbers in ${\mathscr{W}}_{r}(\alpha)$ are at least $\frac{\alpha}{2}$.
Lemma 2.1.
If $12\leq r\leq\alpha^{2}/4+$ 1, then $q>\frac{\alpha}{2}$ for each $q\in{\mathscr{W}}_{r}(\alpha)$.
Proof. Let $\beta:=\frac{\alpha+\sqrt{\alpha^{2}-4r+4}}{2}$. Since $12\leq r\leq\alpha^{2}/4+$ 1, we have $\alpha>\beta\geq\frac{\alpha}{2}$.
We will prove that $q>\beta$ holds for each $q\in{\mathscr{W}}_{r}(\alpha)$. Suppose this conclusion does not hold. By the definition of ${\mathscr{W}}_{r}(\alpha)$ and the fact $\alpha>\beta$, there exists a multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):i\in\llbracket s\rrbracket\}$, where $1\leq s<r$, such that $q_{i}>\beta$ for all $i\in\llbracket s\rrbracket$ and
$q:=\alpha-\sum_{i=1}^{s}q_{i}^{-1}\leq\beta.$ (2)
However, it contradicts the following conclusion:
$q=\alpha-\sum_{i=1}^{s}q_{i}^{-1}>\alpha-(r-1)\beta^{-1}=\beta,$ (3)
where the equality $\alpha-(r-1)\beta^{-1}=\beta$ follows from the assumption $\beta=\frac{\alpha+\sqrt{\alpha^{2}-4r+4}}{2}$. Hence the lemma follows. $\Box$
For example, if $r=2$ and $\alpha=2$, then $r=\frac{\alpha^{2}}{4}+1$. By Lemma 2.1, $q>\frac{2}{2}=1$ for all $q\in{\mathscr{W}}_{2}(2)$. This conclusion coincides with the fact that ${\mathscr{W}}_{2}(2)=\{\frac{i+1}{i}:i\geq 1\}$. By Theorem 1.1, Lemma 2.1 implies that there is no trees $T$ with $12\leq\Delta(T)\leq\lambda(T)^{2}/4+$ 1, which coincides with inequality (1) due to Stevanović [16].
Now we are going to introduce the Perron–Frobenius theorem, by which we develop a result for studying the numbers in ${\mathscr{W}}_{r}(\alpha)$.
The Perron–Frobenius theorem, proved by Oskar Perron [14] in 1907 and Georg Frobenius [9] in 1912, is an important result in the study of eigenvalues and eigenvectors of square matrices. Applying the Perron–Frobenius theorem for irreducible non-negative matrices, the following conclusion for the spectrum radius $\lambda(G)$ of a connected graph $G$ is obtained. Let $\lambda_{i}(G)$ denote the $i$-th largest eigenvalue of $G$. Then, $\lambda(G)=\lambda_{1}(G)$. We say $G$ is non-trivial if $|V(G)|\geq 2$.
@@ 2 The set ${\mathscr{W}}_{r}(\alpha)$
Applying Theorem 2.2, the next conclusion follows directly. Let ${\mathbb{R}}^{+}$ be the set of positive real numbers.
@@
Let $G$ be a connected non-trivial graph and $\alpha$ be a positive number. Then, $\lambda(G)=\alpha$ if and only if there is a mapping $\phi$ from $V(G)$ to ${\mathbb{R}}^{+}$ such that for each $u\in V(G)$,
@@
Now we are going to establish a relation between a mapping $\phi:V(G)\rightarrow{\mathbb{R}}^{+}$ satisfying the condition in (4) and the numbers in ${\mathscr{W}}_{r}(\alpha)$.
@@
Let $T$ be a non-trivial tree and $\alpha=\lambda(T)$. If $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ is a mapping satisfying the condition that for each $u\in V(T)$,
$\alpha\phi(u)=\sum_{w\in N_{T}(u)}\phi(w),$ (5)
then $\frac{\phi(u_{1})}{\phi(u_{2})}\in{\mathscr{W}}_{r}(\alpha)$ for each edge $u_{1}u_{2}$ in $T$, where $r=\Delta(T)$. Therefore, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$.
Proof. Let $u_{1}u_{2}$ be a fixed edge in $T$ and let $T^{\prime}$ be the subtree of $T$ induced by $\{u_{1}\}\cup V(T_{2})$, where $T_{i}$ is the component of $T-u_{1}u_{2}$ (i.e., the subgraph of $T$ obtained by deleting edge $u_{1}u_{2}$) that contains vertex $u_{i}$. Clearly, $u_{1}\in L(T^{\prime})$ (i.e., a leaf in $T^{\prime}$). For convenience, let $T^{\prime}$ denote the directed tree obtained from $T^{\prime}$ such that $u_{1}$ is its unique source (i.e., a vertex with in-degree 00), as shown in Figure 1 (b).
We are now going to prove the following claim by induction. For any vertex $w$ in $T^{\prime}$, let $d_{T^{\prime}}(u_{1},w)$ denote the distance from $u_{1}$ to $w$ in $T^{\prime}$.
Claim 1: For any directed edge $(w_{1},w_{2})$ in $T^{\prime}$, $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$.
Let $d=\max\{d_{T^{\prime}}(u_{1},w):w\in V(T^{\prime})\}$. If $w_{2}$ is a sink of $T^{\prime}$ (i.e., $w_{2}\in L(T)$), then by condition (5), $\alpha\phi(w_{2})=\phi(w_{1})$, implying that $\frac{\phi(w_{1})}{\phi(w_{2})}=\alpha\in{\mathscr{W}}_{r}(\alpha)$. Thus, Claim 1 holds for any directed edge $(w_{1},w_{2})$ of $T^{\prime}$ with $d_{T^{\prime}}(u_{1},w_{2})=d$.
@@
    (a) $T$                                                                  (b)     $T^{\prime}$
Figure 1: Tree $T$ and directed tree $T^{\prime}$ in ${\mathscr{T}}$
Assume that Claim 1 holds for each directed edge $(w^{\prime}_{1},w^{\prime}_{2})$ in $T^{\prime}$ with $d_{T^{\prime}}(u_{1},w^{\prime}_{2})\geq t$, where $2\leq t\leq d$. Now we take any directed edge $(w_{1},w_{2})$ in $T^{\prime}$ such that $w_{2}$ is not a sink of $T^{\prime}$ and $d_{T^{\prime}}(u_{1},w_{2})=t-1$. Let $N_{T}(w_{2})\setminus\{w_{1}\}=\{v_{i}:i\in\llbracket s\rrbracket\},$ where $s=d_{T}(w_{2})-1\leq r-1$. Clearly, $d_{T^{\prime}}(u_{1},v_{i})=d_{T^{\prime}}(u_{1},w_{2})+1=t$ for all $i\in\llbracket s\rrbracket$. By the inductive assumption, $\frac{\phi(w_{2})}{\phi(v_{i})}\in{\mathscr{W}}_{r}(\alpha)$ for all $i\in\llbracket s\rrbracket$. By condition (5),
@@
As $\frac{\phi(w_{2})}{\phi(v_{i})}\in{\mathscr{W}}_{r}(\alpha)$ for all $i\in\llbracket s\rrbracket$, by the definition of ${\mathscr{W}}_{r}(\alpha)$, we have $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$. Hence Claim 1 holds.
It follows from Claim 1 that $\frac{\phi(u_{1})}{\phi(u_{2})}\in{\mathscr{W}}_{r}(\alpha)$ holds for each edge $u_{1}u_{2}$ in $T$. Taking any edge $u_{1}u_{2}\in E(T)$, where $u_{1}\in L(T)$, we have $\alpha\phi(u_{1})=\phi(u_{2})$ by (5), implying that $\alpha^{-1}=\frac{\phi(u_{1})}{\phi(u_{2})}\in{\mathscr{W}}_{r}(\alpha)$. Thus, this result follows. $\Box$
3 Weighted Directed trees
Let ${\mathscr{T}}$ be the set of directed trees $T$ of order at least $2$ which has a unique source. Clearly, for any $T\in{\mathscr{T}}$ with the unique source $z$, the in-degree of each vertex $v\in V(T)\setminus\{z\}$ in $T$ is $1$. For any undirected non-trivial tree and any vertex $z$ in this tree, there is one and only one orientation to obtain a directed tree with $z$ as its unique source.
For any real number $\alpha$ with $\alpha\geq 1$, let ${\mathscr{T}}_{\alpha}$ be the set of trees $T\in{\mathscr{T}}$ such that (8) defines a vertex-weight function $\omega$ from $V(T)$ to $\{0\}\cup{\mathbb{R}}^{+}$:
$\omega(v)=\left\{\begin{array}[]{ll}\alpha,&\mbox{if }od_{T}(v)=0;\\ \alpha-\sum\limits_{w\in V(T)\atop(v,w)\in E(T)}\omega(w)^{-1},&\mbox{otherwise},\end{array}\right.$ (8)
where $od_{T}(v)$ is the out-degree of $v$ in $T$ and $(v,w)\in E(T)$ represents the directed edge from $v$ to $w$ in $T$. For any $T\in{\mathscr{T}}_{\alpha}$, let $\omega_{\alpha,T}$ denote the mapping defined in (8), which is unique with respect to $\alpha$ and $T$. Clearly, $\omega_{\alpha,T}(v)>0$ for all $v\in V(T)$, unless $v$ is the unique source of $T$.
For example, if $T\in{\mathscr{T}}$ is the directed tree obtained from the star $K_{1,3}$ so that its unique source is of out-degree $1$ in $T$, it can be verified that $T$ belongs to both ${\mathscr{T}}_{2}$ and ${\mathscr{T}}_{\sqrt{3}}$, as shown in Figure 2 (a) and (b), but $T\notin{\mathscr{T}}_{\frac{3}{2}}\cup{\mathscr{T}}_{\sqrt{2}}$, as shown in Figure 2 (c) and (d), where the number beside each vertex $v$ in $T$ is the value of $\omega_{\alpha,T}(v)$.
1111111122222222
[…]
$-\frac{9}{2}$$\frac{1}{6}$$\frac{3}{2}$$\frac{3}{2}$
[…]
(a) $T\in{\mathscr{T}}_{2}$                  (b) $T\in{\mathscr{T}}_{\sqrt{3}}$                  (c) $T\notin{\mathscr{T}}_{\frac{3}{2}}$                  (d) $T\notin{\mathscr{T}}_{\sqrt{2}}$
Figure 2: $T\in{\mathscr{T}}_{2}$, $T\in{\mathscr{T}}_{\sqrt{3}}$, but $T\notin{\mathscr{T}}_{\frac{3}{2}}\cup{\mathscr{T}}_{\sqrt{2}}$, where $T$ is oriented from $K_{1,3}$
For non-negative numbers $r$, $\alpha$ and $q$, where $r$ is a positive integer and $\alpha\geq 1$, let ${\mathscr{T}}_{r,\alpha}(q)$ denote the set of $T\in{\mathscr{T}}_{\alpha}$ satisfying the following conditions:
@@ 3 Weighted Directed trees
    $\omega_{\alpha,T}(z)=q$, where $z$ is the unique source of $T$, and

### v2

# An algebraic characterization for spectral radii of trees

@@ Abstract
There have been various results on spectral radii of trees, studying the relations between spectral radius $\lambda(T)$ of a tree $T$ and some other parameters of $T$. In this article, we provide an algebraic characterization for the spectral radii of trees, and apply this new result to study the existence of trees with given spectral radius and maximum degree.
For any positive integer $r$ and positive number $\alpha$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set defined recursively: $\alpha\in{\mathscr{W}}_{r}(\alpha)$, and for any multi-set $\{q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,1\leq i\leq s\}$, if $\beta:=\alpha-\sum\limits_{i=1}^{s}q_{i}^{-1}\geq 0$ and $s\leq r-\left\lceil{\frac{\beta}{\beta+1}}\right\rceil$, then $\beta\in{\mathscr{W}}_{r}(\alpha)$. We first show that for any positive integer $r$ and real number $\alpha\geq 1$, there exists a tree $T$ with maximum degree $\Delta(T)\leq r$ and spectral radius $\lambda(T)=\alpha$ if and only if $0\in{\mathscr{W}}_{r}(\alpha)$. It follows that the set of spectral radii of non-trivial trees is exactly the set of positive numbers $\alpha$ such that $0\in{\mathscr{W}}_{\lfloor\alpha^{2}\rfloor}(\alpha)$. Applying this conclusion, we then prove that for any positive integers $r\geq 2$ and $k$, there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\sqrt{k}$ if and only if $\frac{1}{4}k+1<r\leq k$.
@@
Keywords: tree, maximum degree, characteristic polynomial; eigenvalue; spectral radius
@@ 1 Introduction
All graphs considered in this article are simple and undirected. For a graph $G$, let $V(G)$, $E(G)$ and $\Delta(G)$ denote its vertex set, edge set and maximum degree, respectively. For any $u\in V(G)$, let $N_{G}(u)$ (or simply $N(u)$) denote the set of neighbours of $u$ in $G$, and let $d_{G}(u)$ or (simply $d(u)$) be the size of $N_{G}(u)$, called the degree of $u$ in $G$. Let $L(G)$ denote the set of vertices $v\in V(G)$ with $d_{G}(v)=1$.
Let ${\mathbb{N}}^{+}$ denote the set of positive integers and for any $k\in{\mathbb{N}}^{+}$, let $\llbracket k\rrbracket=\{i\in{\mathbb{N}}^{+}:i\leq k\}$. For a graph $G=(V,E)$, where $V=\{v_{i}:i\in\llbracket k\rrbracket\}$, its adjacency matrix $A(G)$ according to the vertex ordering $v_{1},v_{2},\ldots,v_{n}$ is defined to be the $0-1$ matrix $(a_{i,j})_{n\times n}$, where $a_{i,j}=1$ if and only if $v_{i}v_{j}\in E$. The characteristic polynomial of $G$, denoted by $P(G,\lambda)$, is defined to be the function $det(\lambda I_{n}-A(G))$, where $I_{n}$ is the identical matrix of size $n$. It is known that $P(G,\lambda)$ is independent of the vertex ordering $v_{1},v_{2},\dots,v_{n}$, and the roots of the equation $P(G,\lambda)=0$, called the eigenvalues of $G$, are real numbers (see [6]), and the largest eigenvalue of $G$, denoted by $\lambda(G)$, is called the spectral radius of $G$.
The spectral radii of graphs, which have been studied extensively (see a survey [26] due to Stevanovíc), plays a crucial role in spectral graph theory, with applications in spanning combinatorics, network theory, and even theoretical physics (see [1, 2, 3, 4, 9, 12, 13, 22, 23, 26]). The distribution of spectral radii among different graphs is a topic of ongoing research with connections to various graph properties and invariants (see [7, 12, 13, 16, 20, 22, 23]).
In this article, we focus on the study of spectral radii of trees. There are various known results on the relation between $\lambda(T)$ and other parameters of a tree $T$. Guo and Tan [11] found an upper bound of $\lambda(T)$ in terms of the order and matching number of $T$. Lovász and Pelikán [15] determined the maximal and minimal value of $\lambda(T)$ among all trees $T$ of order $n$, Simić and Tos̆ić [24] found the tree that has the largest spectral radius among all trees with given order and maximum degree, and Stevanović [25] provided a relation between $\lambda(T)$ and $\Delta(T)$: for any tree $T$ with $|V(T)|\geq 3$,
$\frac{1}{4}\lambda(T)^{2}+1<\Delta(T)\leq\lambda(T)^{2}.$ (1)
In this article, we will first provide a characterization on $r$ and $\alpha$ for the existence of trees $T$ with $\Delta(T)=r$ and $\lambda(T)=\alpha$. Let $\Lambda$ denote the set of $\lambda(T)$ for all non-trivial trees $T$ (i.e., trees of order at least $2$). Clearly, $\alpha\geq 1$ for each $\alpha\in\Lambda$. Let ${\mathbb{R}}^{+}$ denote the set of positive real numbers, and for any $r\in{\mathbb{N}}^{+}$ and $\alpha\in{\mathbb{R}}^{+}$, let ${\mathscr{W}}_{r}(\alpha)$ denote the set of non-negative real numbers defined below:
@@
    for any $s\in{\mathbb{N}}^{+}$ and multi-set $\big{\{}q_{i}\in{\mathscr{W}}_{r}(\alpha):q_{i}>0,i\in\llbracket s\rrbracket\big{\}}$, if $\beta:=\alpha-\sum\limits_{1\leq i\leq s}q_{i}^{-1}\geq 0$ and $1\leq s\leq r-\left\lceil{\frac{\beta}{\beta+1}}\right\rceil$, then $\beta\in{\mathscr{W}}_{r}(\alpha)$.
It can be verified that ${\mathscr{W}}_{1}(1)=\{0,1\}$, ${\mathscr{W}}_{1}(\alpha)=\{\alpha\}$ for all $\alpha\in{\mathbb{R}}^{+}$ with $\alpha\neq 1$, ${\mathscr{W}}_{2}(\sqrt{2})=\left\{0,\sqrt{2},(\sqrt{2})^{-1}\right\}$ and ${\mathscr{W}}_{2}(2)=\big{\{}\frac{i+1}{i}:i\geq 1\big{\}}$. Clearly, the property ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}_{r+1}(\alpha)$ holds for any $\alpha\in{\mathbb{R}}^{+}$ and $r\in{\mathbb{N}}^{+}$.
Our first aim in this article is to establish a connection between the existence of a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ and the existence of 00 in the set ${\mathscr{W}}_{r}(\alpha)$.
@@
For any $\alpha\in{\mathbb{R}}^{+}$ and $r\in{\mathbb{N}}^{+}$, there exists a tree $T$ with $\Delta(T)\leq r$ and $\lambda(T)=\alpha$ if and only if $0\in{\mathscr{W}}_{r}(\alpha)$.
A characterization of the set $\Lambda$ follows from Theorem 1.1 and (1) directly.
@@
$\Lambda=\big{\{}\alpha\in{\mathbb{R}}^{+}:0\in{\mathscr{W}}_{\left\lfloor{\alpha^{2}}\right\rfloor}(\alpha)\big{\}}$.
If $r$ is removed from the definition of the set ${\mathscr{W}}_{r}(\alpha)$, we can define the set ${\mathscr{W}}(\alpha)$ as follows:
  1. (i).
    $\alpha\in{\mathscr{W}}(\alpha)$; and
  2. (ii).
    for any multi-set $\big{\{}q_{i}\in{\mathscr{W}}(\alpha):q_{i}>0,i\in\llbracket s\rrbracket\big{\}}$, if $\beta:=\alpha-\sum\limits_{1\leq i\leq s}q_{i}^{-1}\geq 0$, then $\beta\in{\mathscr{W}}(\alpha)$.
Clearly, ${\mathscr{W}}_{r}(\alpha)\subseteq{\mathscr{W}}(\alpha)$ for any $r\in{\mathbb{N}}^{+}$, and the following conclusion holds.
Corollary 1.3.
$\Lambda=\big{\{}\alpha\in{\mathbb{R}}^{+}:0\in{\mathscr{W}}(\alpha)\big{\}}$.
The conclusion of (1) tells that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$ is a necessary condition for the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$. We wonder if this condition is also sufficient.
@@
Our second aim is to apply Theorem 1.1 to prove Conjecture 1.4 for the case $\alpha=\sqrt{k}$ for each $k\in{\mathbb{N}}^{+}\setminus\{1\}$. Certainly, it includes the case $\alpha=k$ for each $k\in{\mathbb{N}}^{+}\setminus\{1\}$.
@@
For any $k,r\in{\mathbb{N}}^{+}\setminus\{1\}$, there exists a tree $T$ with $\lambda(T)=\sqrt{k}$ and $\Delta(T)=r$ if and only if $\frac{1}{4}k+1<r\leq k$.
The remainder of this article is structured as follows. In Sections 2 and 3, we prove Theorem 1.1’s necessity and sufficiency, respectively. In Section 4, we complete the proof of Theorem 1.5. Finally, in Section 5, we provide some open problems arising from our work.
2 The necessity of Theorem 1.1
The Perron–Frobenius theorem, proved by Oskar Perron [19] in 1907 and Georg Frobenius [10] in 1912, is an important result in the study of eigenvalues and eigenvectors of square matrices. Applying the Perron–Frobenius theorem for irreducible non-negative matrices, the following conclusion for the spectral radius $\lambda(G)$ of a connected graph $G$ is obtained. Let $\lambda_{i}(G)$ denote the $i$-th largest eigenvalue of $G$. Then, $\lambda(G)=\lambda_{1}(G)$. We say $G$ is non-trivial if $|V(G)|\geq 2$.
@@
Applying Theorem 2.1, the next conclusion follows directly.
@@
Let $G$ be a connected non-trivial graph and $\alpha$ be a positive number. Then, $\lambda(G)=\alpha$ if and only if there is a mapping $\phi:V(G)\mapsto{\mathbb{R}}^{+}$ such that for each $u\in V(G)$,
@@
Note that for any mapping $\phi:V(G)\mapsto{\mathbb{R}}^{+}$, the condition of (2) holds if and only if $\phi$ is an eigenvector of $A(G)$ with respect to $\lambda(G)$. Now we are going to prove the following lemma by which the necessity of Theorem 1.1 follows directly.
@@
Let $T$ be a non-trivial tree and $\alpha=\lambda(T)$, and let $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ be an eigenvector of $A(T)$ with respect to $\alpha$. Then both $\frac{\phi(u_{1})}{\phi(u_{2})}$ and $\frac{\phi(u_{2})}{\phi(u_{1})}$ belong to ${\mathscr{W}}_{r}(\alpha)$ for each $u_{1}u_{2}\in E(T)$, where $r=\Delta(T)$.
Proof. Since $\phi:V(T)\mapsto{\mathbb{R}}^{+}$ is an eigenvector of $A(T)$ with respect to $\alpha=\lambda(T)$, (2) holds for each $u\in V(T)$. Let’s first prove the following claim.
Claim 1: For an arbitrary vertex $z$ in $T$, if $T^{\prime}$ is the orientation of $T$ with $z$ as its source, as shown in Figure 1, then $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$ holds for every directed edge $(w_{1},w_{2})$ in $T^{\prime}$.
@@
Figure 1: A directed tree $T^{\prime}$ with $z$ as its unique source
We will prove Claim 1 by induction. For any vertex $w$ in $T^{\prime}$, let $d_{T^{\prime}}(z,w)$ denote the distance from $z$ to $w$ in $T^{\prime}$, and let $d=\max\{d_{T^{\prime}}(z,w):w\in V(T^{\prime})\}$.
If $w_{2}$ is a sink of $T^{\prime}$, then $w_{2}$ is a leaf of $T$, and the condition of (2) yields that $\alpha\phi(w_{2})=\phi(w_{1})$, implying that $\frac{\phi(w_{1})}{\phi(w_{2})}=\alpha\in{\mathscr{W}}_{r}(\alpha)$. Thus, Claim 1 holds whenever $w_{2}$ is a sink of $T^{\prime}$. Clearly, if $d_{T^{\prime}}(u_{1},w_{2})=d$, then $w_{2}$ is a sink of $T^{\prime}$ and thus Claim 1 also holds.
Assume that Claim 1 holds for each directed edge $(w^{\prime}_{1},w^{\prime}_{2})$ in $T^{\prime}$ with $d_{T^{\prime}}(z,w^{\prime}_{2})\geq t$, where $2\leq t\leq d$. Now we take any directed edge $(w_{1},w_{2})$ in $T^{\prime}$ such that $w_{2}$ is not a sink of $T^{\prime}$ and $d_{T^{\prime}}(z,w_{2})=t-1$. Assume that $N_{T}(w_{2})\setminus\{w_{1}\}=\{v_{i}:i\in\llbracket s\rrbracket\},$ where $s=d_{T}(w_{2})-1\leq r-1$. Clearly, for each $i\in\llbracket s\rrbracket$, $(w_{2},v_{i})$ is a directed edge in $T^{\prime}$ and $d_{T^{\prime}}(z,v_{i})=d_{T^{\prime}}(z,w_{2})+1=t$. By the inductive assumption, $\frac{\phi(w_{2})}{\phi(v_{i})}\in{\mathscr{W}}_{r}(\alpha)$ for all $i\in\llbracket s\rrbracket$. By the condition of (2),
@@
As $s<r$, by the definition of ${\mathscr{W}}_{r}(\alpha)$, we have $\frac{\phi(w_{1})}{\phi(w_{2})}\in{\mathscr{W}}_{r}(\alpha)$. Hence Claim 1 holds.
Since vertex $z$ in Claim 1 is an arbitrary vertex in $T$, Claim 1 implies that both $\frac{\phi(u_{1})}{\phi(u_{2})}$ and $\frac{\phi(u_{2})}{\phi(u_{1})}$ belong to ${\mathscr{W}}_{r}(\alpha)$ for each $u_{1}u_{2}\in E(T)$. $\Box$
Proof of the necessity of Theorem 1.1. We follow the conditions in Lemma 2.3. Take any edge $u_{1}u_{2}\in E(T)$, where $u_{1}$ is a leaf of $T$. We have $\alpha\phi(u_{1})=\phi(u_{2})$ by (2), implying that $\alpha^{-1}=\frac{\phi(u_{2})}{\phi(u_{1})}\in{\mathscr{W}}_{r}(\alpha)$ by Lemma 2.3. Thus, $0=\alpha-\frac{1}{\alpha^{-1}}\in{\mathscr{W}}_{r}(\alpha)$ by definition.
Hence the result holds. $\Box$
3 The sufficiency of Theorem 1.1
In this section, we always assume that $r\in{\mathbb{N}}^{+}$ and $\alpha\in{\mathbb{R}}^{+}$. Let ${\mathscr{T}}_{r,\alpha}$ denote the set of ordered pairs $(T,\omega)$, where $T$ is a directed tree with a unique source $z$ and $\omega$ is a mapping from $V(T)$ to ${\mathscr{W}}_{r}(\alpha)$ such that for each $v\in V(T)$,
@@
    $od_{T}(v)\leq r$ if $v=z$ and $\omega(z)=0$, and $od_{T}(v)\leq r-1$ otherwise; and